OpenStudy (anonymous):

The officejet printer can copy Sue's dissertation in 10 minutes. The laserjet printer can copy the same document in 22 minutes. If the two machines work together how long would they take to copy the dissertation?

OpenStudy (saifoo.khan):

Nice Photo!

OpenStudy (anonymous):

Thanks.

OpenStudy (saifoo.khan):

1/10+1/22=1/x (10+22)/10*22=1/x 32/220=1/x 32x=220 x=220/32 x=6.875 minutes

OpenStudy (anonymous):

hey saifooo

OpenStudy (saifoo.khan):

Hi Honey..

OpenStudy (anonymous):

saif?

OpenStudy (saifoo.khan):

Rachel?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[x = 55/8\]